Circular No. 6794

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997eg IN NGC 5012 A. V. Filippenko and A. J. Barth, University of California at Berkeley,

report: “A spectrum (range 430–680 nm, resolution 0.5 nm) obtained on Dec. 20 with the Keck-2 telescope reveals that SN 1997eg (cf. IAUC 6790) is a peculiar type-II supernova. The spectrum contains strong emission lines superposed on a featureless blue continuum. Although generically referred to as ‘type IIn’ supernovae, this object is most similar to the subset whose prototype is SN 1988Z (e.g., MNRAS 250, 786, and 262, 128). Hα has three strong components: unresolved (FWHM < 200 km/s), intermediate (FWHM ∼ 2000 km/s), and broad (FWHM ∼ 12 000 km/s). The broad He I 587.6-nm emission line is nearly as strong as the corresponding component of Hα, suggesting a very high helium abundance. The unresolved lines come from a very dense circumstellar medium (n > ∼ 10^{8} cm^{−3}): the [O III] 436.3nm emission is roughly as strong as [O III] at 500.7 nm. Moreover, they are characterized by very high excitation: He II 468.6-nm and [Fe X] 637.5-nm are prominent. Based on the behavior of previous examples of this rare subset of type-IIn supernovae, we predict that SN 1997eg will become very luminous at x-ray and radio wavelengths due to strong interaction between the ejecta and circumstellar material.”

GRB 971214 E. Waxman, Institute for Advanced Study, Princeton, writes: “The

data of the detected optical and x-ray afterglow from GRB 971214 (IAUC 6787–6791 and references therein, plus private comm. by L. Piro et al.) were used to make approximate predictions for its behavior at different wavelengths, which may be useful in planning observations. The detected flux reported in the first two days of observations is consistent with general fireball-model predictions, yielding optical-band magnitudes predicted by m_{o} + 3 log t, where t is measured in days and m_{o} = 24.6 for B, 23.4 for V, 22.9 for R, 22.1 for I, and 21.1 for J. The x-ray flux is predicted at ∼ 4 keV by 5 × 10^{−13} t^{−1.2} erg s^{−1} cm^{−2}. Simultaneous observations at different optical colors (and at x-rays) would provide strong constraints on the model. There is some indication for extinction at the source (incorporated in the above numbers); if real, it implies that absorption lines should be detected (EW ∼ 0.004 nm for Mg II), with the equivalent width significantly changing on timescales of ∼ 1 day — and spectra taken at different times would therefore be useful. Monitoring of optical fluxes on < 1-day timescales would also be useful in constraining fireball characteristics.”

1997 December 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green